1. Choose the equation of the parabola with a vertex at the origin and a focus at (0, –3).

1: Y=-1/3X
2: Y=1/12X^2
3: Y=-1/3X^2
4: Y=-1/12X^2

Answer by Alan3354(21615)

1. Choose the equation of the parabola with a vertex at the origin and a focus at (0, –3).
1: Y=-1/3X
2: Y=1/12X^2
3: Y=-1/3X^2
4: Y=-1/12X^2
----------------
It's 4, y = -x^2/12